OpenStudy (anonymous):

Given the equation y=3x, find average rate of change of y from x1=0 to x2=8

OpenStudy (anonymous):

I know the avg rate of change is (y2-y1)/(x2-x1)

OpenStudy (anonymous):

i got y1=0 and y2=64

OpenStudy (anonymous):

and as far as 64/8

OpenStudy (anonymous):

I have a few problems like this so I dont know what to do for them

jimthompson5910 (jim_thompson5910):

how did you get y2 = 64?

jimthompson5910 (jim_thompson5910):

I'm getting y = 3x y = 3*8 y = 24 so y2 = 24 when x2 = 8

OpenStudy (anonymous):

oh that makes more sense. webwork's examples are confusing

jimthompson5910 (jim_thompson5910):

hmm, well anyways, what new final answer do you get?

OpenStudy (anonymous):

WebWork said to do it as (x2)^2 so I squared 8. I got 3

jimthompson5910 (jim_thompson5910):

isn't the original equation y = 3x and not y = x^2 ?

OpenStudy (anonymous):

yea

jimthompson5910 (jim_thompson5910):

just a bad typo it seems

jimthompson5910 (jim_thompson5910):

yeah the avg rate of change is 3

jimthompson5910 (jim_thompson5910):

for any linear equation, the avg rate of change is the slope of the line

jimthompson5910 (jim_thompson5910):

you'll see that avg rate of change visually corresponds to the slope of the secant line